OpenStudy (anonymous):

use the method of variation of parameters to find a particular solution of the given equation x^2y''-4xy'+6y=x^7 where y1 = x^2 and y2 = x^3 use the method of variation of parameters to find a particular solution of the given equation x^2y''-4xy'+6y=x^7 where y1 = x^2 and y2 = x^3 @Mathematics

OpenStudy (anonymous):

i get y = x^6/12 + c1x^3 - c2x^2

OpenStudy (jamesj):

x^6/12 isn't right.

OpenStudy (jamesj):

Let the particular solution be yp = Ax^7. Then yp ' = 7Ax^6, yp'' = 42Ax^5 thus x^2y''-4xy'+6y = (42A - 28A + 6A) x^7 = x^7 and therefore A = 1/20